# A Miyaoka-Yau inequality for hyperplane arrangements in $\mathbb{CP}^{n}$

### v2

@@ 7.3 Complex reflection arrangements
Consider the subgroup $G_{L}\subset G$ made of elements that fix $L$. By [OT92, Theorem 6.25], the group $G_{L}$ generated by the reflections in $G$ whose reflecting hyperplanes belong to the localized arrangement $\mathcal{H}_{L}$. We identify $L^{\perp}=\mathbb{C}^{p}$ with $p=\operatorname{codim}L$ and $G_{L}\subset U(p)$, as the group $G_{L}$ acts faithfully on $L^{\perp}$. The reflecting hyperplanes of $G_{L}$ are $H\cap L^{\perp}$ for $H\in\mathcal{H}_{L}$. Since, by assumption, the arrangement $\mathcal{H}_{L}$ is irreducible, the action of $G_{L}$ on $\mathbb{C}^{p}$ is also irreducible. Equation (7.26) follows by Lemma 7.22 . ∎
@@ A.1.2 Tuples of filtrations and splittings
The filtrations $\mathcal{F}^{1},\ldots,\mathcal{F}^{k}$ are compatible if $\mathcal{F}$ admits a splitting.
@@ A.1.5 Nested sets of projective subspaces
A nested set in $\mathbb{CP}^{n}$ is a finite set $\mathcal{S}$ of non-empty and proper projective subspaces $L\subsetneq\mathbb{CP}^{n}$ such that, for any collection $L_{1},\ldots,L_{k}\in\mathcal{S}$ of pairwise non comparable elements (i.e. $L_{i}\not\subset L_{j}$ for $i\neq j$), the following holds:
@@ A.4.2 Multivector fields tangent to hyperplanes
We are left to show that $\operatorname{Tan}(\mathbf{v})\subset\operatorname{Tan}(\mathbf{v}^{\prime})$. Notice that, since $\mathbf{v}^{\prime}\in T_{1}$, the coordinate hyperplanes are contained in $\operatorname{Tan}(\mathbf{v}^{\prime})$. Now, let $H$ be a hyperplane in $\operatorname{Tan}(\mathbf{v})$ different from the coordinate hyperplanes and take a point $x\in H$ outside the origin. The multivector $\mathbf{v}^{\prime}_{x}$ is obtained by successive contractions of $\mathbf{v}_{x}$ by elements in $(T_{x}\mathbb{C}^{n+1})^{*}$. Since $\mathbf{v}_{x}$ is tangent to $H$, it follows from Lemma A.45 that $\mathbf{v}^{\prime}_{x}$ is tangent to $H$. Therefore, $H\in\operatorname{Tan}(\mathbf{v}^{\prime})$ and the proposition follows. ∎
@@ Appendix C Miyaoka-Yau and Kähler-Einstein metrics
where $(X^{n},\omega_{\operatorname{KE}})$ is a compact Kähler-Einstein manifold, $\mathring{\textup{Riem}}$ is the trace free part of the Riemann curvature tensor of $\omega_{\operatorname{KE}}$, and $\lambda_{n}$ is a positive dimensional constant. As a consequence, if equality holds in the Miyaoka-Yau inequality, then $\omega_{\operatorname{KE}}$ must have constant holomorphic sectional curvature.

### v3

@@ 7.3 Complex reflection arrangements
Consider the subgroup $G_{L}\subset G$ made of elements that fix $L$. By [OT92, Theorem 6.25], the group $G_{L}$ is generated by the reflections in $G$ whose reflecting hyperplanes belong to the localized arrangement $\mathcal{H}_{L}$. We identify $L^{\perp}=\mathbb{C}^{p}$ with $p=\operatorname{codim}L$ and $G_{L}\subset U(p)$, as the group $G_{L}$ acts faithfully on $L^{\perp}$. Thus $\mathcal{H}_{L}$ is a complex reflection arrangement. Since, by assumption, the arrangement $\mathcal{H}_{L}$ is irreducible, the action of $G_{L}$ on $\mathbb{C}^{p}$ is also irreducible. Equation (7.26) follows by Lemma 7.22 . ∎
@@ A.1.2 Tuples of filtrations and splittings
The filtrations $\mathcal{F}^{1},\ldots,\mathcal{F}^{k}$ are compatible if the tuple $\mathcal{F}=(\mathcal{F}^{1},\ldots,\mathcal{F}^{k})$ admits a splitting.
@@ A.1.5 Nested sets of projective subspaces
A nested set in $\mathbb{CP}^{n}$ is a finite set $\mathcal{S}$ of non-empty and proper projective subspaces $L\subsetneq\mathbb{CP}^{n}$ such that, for any collection $L_{1},\ldots,L_{k}\in\mathcal{S}$ with $k\geq 2$ of pairwise non comparable elements (i.e. $L_{i}\not\subset L_{j}$ for $i\neq j$), the following holds:
@@ A.4.2 Multivector fields tangent to hyperplanes
We are left to show that $\operatorname{Tan}(\mathbf{v})\subset\operatorname{Tan}(\mathbf{v}^{\prime})$. Notice that, since $\mathbf{v}^{\prime}\in T_{1}$, the coordinate hyperplanes are contained in $\operatorname{Tan}(\mathbf{v}^{\prime})$. Now, let $H$ be a hyperplane in $\operatorname{Tan}(\mathbf{v})$ different from the coordinate hyperplanes and take a point $x\in H$ outside the origin and not lying in any of the coordinate hyperplanes. The multivector $\mathbf{v}^{\prime}_{x}$ is obtained by successive contractions of $\mathbf{v}_{x}$ by elements in $(T_{x}\mathbb{C}^{n+1})^{*}$. Since $\mathbf{v}_{x}$ is tangent to $H$, it follows from Lemma A.45 that $\mathbf{v}^{\prime}_{x}$ is tangent to $H$. Therefore, $H\in\operatorname{Tan}(\mathbf{v}^{\prime})$ and the proposition follows. ∎
@@ Appendix C Miyaoka-Yau and Kähler-Einstein metrics
where $(X^{n},\omega_{\operatorname{KE}})$ is a compact Kähler-Einstein manifold, $\mathring{\textup{Riem}}$ is the trace free part of the Riemann curvature tensor of $\omega_{\operatorname{KE}}$, and $\lambda_{n}$ is a positive constant depedning on dimension. As a consequence, if equality holds in the Miyaoka-Yau inequality, then $\omega_{\operatorname{KE}}$ must have constant holomorphic sectional curvature.